o.elor.eom/HM.tpx2 -0.669277750288478659 1547785002293 Below are nominal GDP and GDP deflators for four years Nominal GDP in billions GDP 100 103 105 106 Year 2 Year 4 24,280 a. Calculate real GDP in each year. Instructions: Round your responses to the nearest whole number For Year 1, GDP is $ For Year 2, GDP is S For Year 3, GDP is S For Year 4, GDP is $ billion. billion. b. Did the percentage change in nominal GDP exceed the percentage change in rsal GDP in any of the last three years listed? Year 2: Year 3: (Click to select) a. c. In which year did societys welfare increase the most? (Click to select)
Answer #1

GDP deflator=nominal GDP/Real GDP

Real GDP=Nominal GDP/GDP deflator

Year 1 real GDP=20000/100=200

Year 2 real GDP=20560/103=199.61165

Year 3 real GDP=21495/104=206.682692

Year 4 real GDP=22280/106=210.188679

2.In year 2

%change in nominal gdp=(20560-20000/20000)*100=2.8%.

% change in real gdp=(199.61165-200/200)*100=−0.195%

Answer-yes

In year 3

%change in nominal gdp=(21495-20560/20560)*100=4.54%

%change in real gdp=(206.68-199.61/199.61)*100=3.54%

Answer-yes

In year 3

%change in nominal gdp=(22280-21495/21495)*100=3.65%

%change in real gdp=(210.18-206.68/206.68)=1.69%

Answer-yes.

Maximum welfare will be in the year of highest change in real GDP=year 3